OpenStudy (anonymous):

if cos t = -2/5 what does cos(2pi-t) ?

OpenStudy (sidsiddhartha):

cos(2pi-t)=-cost=2/5

OpenStudy (anonymous):

I thought that too, but apparently it equals -2/5 in my book -_-

OpenStudy (anonymous):

@myininaya do you know ?

OpenStudy (anonymous):

@amistre64 ?

OpenStudy (anonymous):

smfh i think my book got it wrong ? =/

myininaya (myininaya):

\[\cos(2\pi-t)=\cos(2\pi)\cos(t)+\sin(2\pi)\sin(t)=1 \cos(t)+0\sin(t)=1\cos(t)=\cos(t)\]

myininaya (myininaya):

book isn't wrong in this case